OpenStudy (anonymous):

find the integers,given the following sum is-3 product is -10

OpenStudy (anonymous):

how 2 do this?

OpenStudy (anonymous):

2; -5

OpenStudy (anonymous):

a+b = -3 ab =-10 a=-3-b (-3-b)b=-10 \[b ^{2}+3b-10=0\] b=2 or -5